at cruising speed, and airplance flew 500 kilometers with the wind in the same amount of time it took on fly 400 kilometers against the wind. if the constant speed of the wind is 20kph, what is the cruising speed of the airplane
Answer by stanbon(57984)

at cruising speed, an airplane flew 500 kilometers with the wind in the same amount of time it took to fly 400 kilometers against the wind. if the constant speed of the wind is 20 kph, what is the cruising speed of the airplane
---------------------
With-wind DATA:
distance = 500 km ; rate = p+20 kph ; time = 500/(p+20) hrs.
-------------------
Against-wind DATA:
distance = 400 km ; rate = p-20 kph ; time = 400/(p-20) hrs.
-------------------
Equation:
time = time
500/(p+20) = 400/(p-20)
---
5/(p+20) = 4/(p-20)
Cross-multiply:
5(p-20) = 4(p+20)
5p-100 = 4p+80
p = 180 kph (speed of the plane in still air)
===================
cheers,
stan H.
===================